OpenStudy (anonymous):

solve 2w+3 ___________ =1 -3

OpenStudy (anonymous):

2w+3 ------- =2w+3\[\div\] -3 is equal to one -3

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

Isolate w to get (2w+3)/(-3) = 1 2w+3 = 1*(-3) 2w+3 = -3 2w = -3-3 2w = -6 w = -6/2 w = -3